Melvin Myers, a traveling salesman, has company medical insurance as well as travel insurance. The company he works for pays 80 percent of the medical insurance and 75 percent of the travel insurance. If the medical insurance has an annual cost of $3,500 and the travel insurance costs $1,200 per year, what total amount is deducted from his weekly paycheck for insurance?

a $9.04 b $36.15 c $19.23 d $29.23
Answer #1

Annual cost of medical insurance = $3,500

Company pays for 80% of the medical insurance. Thus, Melvin pays for the remaining 20% of the medical insurance.

Melvin’s annual contribution for medical insurance = $3,500 * 20% = $700

Melvin’s weekly contribution for medical insurance = $700 / 52 Weeks = $13.46 (Rounded Off)

Annual cost of travel insurance = $1,200

Company pays for 75% of the travel insurance. Thus, Melvin pays for remaining 25% of the travel insurance.

Melvin’s annual contribution for travel insurance = $1,200 * 25% = $300

Melvin’s weekly contribution for travel insurance = $300 / 52 Weeks = $5.77 (Rounded Off)

Melvin’s deduction from weekly paycheck for insurance = Weekly contribution for medical insurance + weekly contribution for travel insurance = $13.46 + $5.77 = $19.23 (Option c)